You are considering the purchase of a Pure Discount Bond with a Face Value of $1,000, which matures in seven years. In the markets this bond is selling for $771.27. If you purchase the bond at this price what is the Yield-to-Maturity (YTM) on the investment? (The answer is a percent, round your answer to two decimal places, e.g. 4.75)
Answer #1

rate =r = (fv/pv)^(1/t)-1 yield to maturity =r= (if annual compounding 3.78% ((1000/771.27)^(1/7)-1) rate =r = (fv/pv)^(1/t)-